OpenStudy (anonymous):

The area of a rectangular garden is given by the trinomial x2 + 6x – 27. What are the possible dimensions of the rectangle? Use factoring.

hero (hero):

A = x^2 + 6x - 27 A = x^2 + 9x - 3x - 27 A = x(x+9)-3(x+9) A = (x+9)(x-3) A = lw l = (x+9) w =(x-3)

OpenStudy (batman31):

True, but your math is off. It would actually be L=(x+10), W=(x-3). but you r process was correct.